OpenStudy (anonymous):

Using a sheet of graph paper, solve the following system of equations graphically. Be sure to show any work, use a straight edge, and be neat. When you are finished, enter the solution set in the box below and turn in your graph paper to your teacher. x + y = 0 x - y + 2 = 0

OpenStudy (anonymous):



OpenStudy (anonymous):

omg i love you!! thanks :'D

OpenStudy (anonymous):

You can check it by doing: \[x+y=x-y+2\]\[2y=2\]\[y=1\]\[x+1=o\]\[x=-1\]\[(-1,1)\]